OpenStudy (anonymous):

find the equation of the line useing the following info show all your work (-8,-8),m= 3/4

OpenStudy (anonymous):

(-8,-8),m=3/4 (y-ysub1)=m(x-xsub1) (y-(-8))=3/4(x-(-8)) y+8=3/4(x+8) 4y+32=3x+24 3x-4y=8